xx EOSHybrid

EOSxHybrid

Harry Dimmelmeier, Ian Hawke, Christian Ott

Date

Abstract

EOSxHybrid. The equation of state used for “simple” core collapse simulations.

1 The equations This equation provides the hybrid polytropic / ideal gas equation of state used by Dimmelmeier et al.x[1] for supernova collapse. A thorn wanting to use this needs to use the CactusEOS interface found in EOSxBase. The equations are

where These expressions do not include the conversions between cgs and Cactus units that are necessary inside the code.

For more details, a recent paper to start with would bex[1].

References

[1]xxxH. Dimmelmeier, J. Novak, A. Font, J.xM. Ibanez, E. Mueller Combining spectral and shock-capturing methods: A new numerical approach for 3D relativistic core collapse simulations Phys. Rev. D71 064023 (2005)

2 Parameters

11eosxgammaxsupernuclear 10Scope: restrictedx01  REAL

11Description: Gammax2 for EOS

11  Range  10  x01  Default: 1.66666666666666

11:  10

11

11eosxgammaxth  10Scope: restrictedx01  REAL

11Description: Gammaxth for EOS

11  Range  10  x01  Default: 1.5

11:  10

11

11rhoxnuc  10Scope: restrictedx01  REAL

11Description: Nuclear matter density

11  Range  10  x01  Default: 1.e-10

11:  10

11

11eosxgamma  10Scope: shared from EOSx2DxPOLYTROPEx01REAL

11

11eosxk  10Scope: shared from EOSx2DxPOLYTROPEx01REAL

11gammaxini  10Scope: shared from EOSx2DxPOLYTROPEx01REAL

11

3 Interfaces

General Implements: eosxhybrid

Inherits: hydrobase eosx2dxpolytrope

Grid Variables

3.0.1 PRIVATE GROUPS

11 x Group Names x 10x Variable Names xx01 Details x11x
11pressurexcontributions10 x01 compact110
11                   10 pressurexpoly x01 description11Polytropic and thermal contributions to the pressure
11 10 pressurexth x01 dimensions113
11                   10  x01 distribution11DEFAULT
11                   10  x01 group type11GF
11 10 x01 tags11Prolongation=”None”
11                   10  x01 timelevels111
11                   10  x01variable type11REAL

11

4 Schedule This section lists all the variables which are assigned storage by thorn EinsteinEOS/EOSxHybrid. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxSTARTUP x eosxhybridxstartup x setup the eos used by dimmelmeier et al. in core collapse x

11x11After: 11eosxpolytropexstartup
11 x 11Language:11fortran
11 x 11Type: 11function

11

CCTKxANALYSIS x checkxpolyxcontrib x output the polytropic and thermal contributions x

11x11Language:11fortran
11 x 11Storage: 11pressurexcontributions
11 x 11Triggers: 11pressurexcontributions
11 x 11Type: 11function